OpenStudy (anonymous):

MEDAL GIVEN b^2+8b+15 divided by b+5

OpenStudy (bibby):

Factor the top half. Hint 15 has the factors 3,5

OpenStudy (anonymous):

what does that mean? Factor the top half?

OpenStudy (bibby):

Simplify the top bit\[\frac{b^2+8b+15 }{ b+5 } = \frac{ (b+?)(b+?) }{ b+5 }\]

OpenStudy (anonymous):

I haven't learned how to do that yet... can you explain how you find what's in the ?

OpenStudy (bibby):

you want to get the quadratic into it's factored form. since it's in the form of b^2 + ?b + ? you start by listing the factors of 15 (1,15) and (3,5) now you need to pick from these The numbers should combine into 8 in some way. here (b+?)(b+?) go look up a review site on how to factor quadratics

OpenStudy (anonymous):

okay

OpenStudy (bibby):

Which part should I clear up? That was a terrible explanation of factoring quadratics but you need to know how to do this before you can answer

OpenStudy (anonymous):

I just haven't learned anything about factoring quadratics and this is on my chapter test

OpenStudy (bibby):

k, I guess I can point out that you should be able to match patterns. If they're supposed to cancel out, you can be sure that there will be a b/5 in the numerator.\[\frac{ (b+5)(b+?) }{ (b+5) }\] can you guess what the other number is? [(1,15)(3,5)]

OpenStudy (anonymous):

15?

OpenStudy (bibby):

nope. The 3 and 5 go together. Can you simplify this?\[\frac{ (b+5)(b+3) }{ (b+5) }\]

OpenStudy (anonymous):

would i multiply b+5 and b+3

OpenStudy (bibby):

that would just give you the thing you started with. Factoring means to break up into the "Facotrs" that make the thing up. You can cancel the b+5s out so that you're left with b+3